1. A man is now 42 years old and his friend is 33 years old. How many years ago was the man twice as old as his friend was then?
2. Paul is 3 years younger than his friend Peter. In seven years the product of their ages will be five more than the product of their ages 5 years ago. How old are they now?
please answer this.. I really need this

Found 2 solutions by ankor@dixie-net.com, Mathtut:
Answer by ankor@dixie-net.com(15647)

I want to do more than answer it, I want to show you how to do it.
:
A man is now 42 years old and his friend is 33 years old. How many years ago was the man twice as old as his friend was then?
:
Let t = number of yrs for this to be true
:
Subtract t from both sides and make an equation:
:
Man, t yrs ago, was twice as old, as his friend t yrs ago
42 - t = 2(33 - t)
42 - t = 66 - 2t; multiply what's inside the brackets
Some basic algebra operations
+2t - t = 66 - 42
t = 24 yrs ago
:
We can prove this
42 - 24 = 18;
33 - 24 = 9
:
:
2. Paul is 3 years younger than his friend Peter. In seven years the product of their ages will be five more than the product of their ages 5 years ago. How old are they now?
:
This problem is impossible, there is no way that the product 7 yrs from now
will be only 5 more than the product 5 yrs ago




Answer by Mathtut(3670)

1. They kind of through a curve ball on this problem... lets call the number of years ago x. and the man we will call A and the friend B. so this is how we write the equation (A-x)=2(B-x) we know that A=42 and B=33 so plug in
42-x=2(33-x) simplify 42-x=66-2x x=24 so the answer is 24 years ago
2.Another curve... something seems to be wrong in the second question: to multiply two ages together in seven years from today as opposed to muliplying those ages 5 years ago and those products only have a difference of 5 is not possible imo